A ball with a mass of 0.30 kg hits the wall at a speed of 5.0 m/s and bounces back at a speed of 4.0 m/s.


Calculate the change in ball momentum.


1
Expert's answer
2021-01-21T18:12:57-0500
"\Delta p=m(v_f-v_i)=0.30\ kg\cdot(-4.0\ \dfrac{m}{s}-5.0\ \dfrac{m}{s})=-2.7\ \dfrac{kgm}{s}."


The sign minus means that the change in ball momentum directed away from the wall.

Answer:

"\Delta p=2.7\ \dfrac{kgm}{s}," away from the wall.